OpenStudy (anonymous):

8. Test the series for convergence or divergence: summantion (n= infinite to1)[n!/(2.5.8..............(3n+2)

OpenStudy (zarkon):

ratio test

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}(n!/2.5.8..................(3n+2))\] its like this

OpenStudy (zarkon):

ok...use the ratio test

OpenStudy (anonymous):

but how?

OpenStudy (zarkon):

compute \[\lim_{n\rightarrow\infty}\left|\frac{a_{n+1}}{a_n}\right|\]

OpenStudy (anonymous):

can you show me all steps please :(

OpenStudy (zarkon):

all of them...hmmm...I'm pretty lazy...I'll just get you started

OpenStudy (zarkon):

\[a_n=\frac{n!}{2\cdot5\cdot8\cdots(3n+2)}\]

OpenStudy (zarkon):

\[\frac{a_n+1}{a_n}=\frac{\frac{(n+1)!}{2\cdot5\cdot8\cdots(3(n+1)+2)}}{\frac{n!}{2\cdot5\cdot8\cdots(3n+2)}}\]

OpenStudy (zarkon):

\[=\frac{n+1}{3(n+1)+2}\]

OpenStudy (zarkon):

take limit

OpenStudy (anonymous):

love you for this :))

OpenStudy (anonymous):

next ?

OpenStudy (zarkon):

did you find the limt